OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m =-1/2 and the point (-2, -3).

OpenStudy (anonymous):

y(x)=mx+b=-x/2+b y(-2)=-3<=> -(-2)/2+b=-3<=>b=-4 y(x)=-x/2-4

OpenStudy (anonymous):

standard form is y = mx + b so -3 = -1/2(-2) + b -3 = 1 + b -3 - 1 = b -4 = b y = -2x - 4

OpenStudy (amistre64):

ax+by=c is standard form; where slope = -a/b x + 2y = c ;calibrate with the point given to find the value for "C"

OpenStudy (anonymous):

Standard form is Ax + By = c. Right now, we have y + 3 = (-1/2)(x + 2), or y = (-1/2)x - 4 That is 2y = -2x - 8. Add 2x. 2x + 2y = -8

OpenStudy (anonymous):

sorry in the final ans i made a slight mistake, it is...... 1 y = - ---- x - 4 2

OpenStudy (amistre64):

you people are insane... you know that right :)